A 10 kg mass hangs on a rope wrapped around a freely rotating 2.00 kg cylinder with a radius of 10.0 cm. Determine the tension in the rope. Use g = 9.8 m/s^2 and I = 1/2 mr^2
Answer #1

T= m(g-a)
T= 10*(9.8-a) -------1

T*r = I*a/r
T*0.1 = 0.5x2x0.12xa/0.1
T= a -------------2

T= 10(9.8-T)

T= 8.91 N